Status of MICE Muon Beamline¶
- Status of MICE Muon Beamline
This list is for tracking the state between BLOCs, and should be viewed before an expert is called. This is complementary to the whiteboard in the control room.
Access / Construction Work¶
- Hall is open.
- PRY construction continues - hence crane activity
Interlocks and Controls¶
- Target gate valve is open
- Beamstop is closed
Services (electricity / water / air / air-conditioning / network)¶
- No service interuptions scheduled.
- Airconditioning unit 1 is isolated, the others are operational.
- The beamline polarity is set to positive polarity
- A new control board has been installed in Q5 MPS to solve ramping issue.
- The Linde Fridge is cooling down after being serviced
- The PSU has undergone a full-current stability test
- If it trips on overtemperature, leave as-is (do NOT reset interlocks) so that the specific channel can be identified
- Proton Absorber:
- need to further close damping valves
- Powerpack 2 has a cracked hose fitting awaiting replacement and is out of use
- No issues